# Eigenvalue Tracking of Large-Scale Systems Impacted by Time Delays

### v1

@@ III-D Communication Delays with Noise and Data Dropouts
The derivations above consider constant delays. In this section, we discuss how they can be conveniently generalized to capture the effects of realistic wide area measurement system (WAMS) latencies with noise and data packet dropouts. To this end, we consider the composite time-varying delay model proposed in [1], wherein for the purpose of SSSA, packet dropouts and noise are captured through the functions $h_{p}(s)$ and $h_{s}(s)$, respectively:
@@
where $\alpha$ and $b$ are the scale and shape factor of the Gamma distribution; $p_{dr}$ is the packet dropout rate; and $T$ is the normal delivery period for each data packet.
@@
where $\tau_{0}$ is the constant component of the WAMS delay. By setting $\boldsymbol{\rm S}_{T}=h_{p}(s)h_{s}(s)\boldsymbol{\rm A}_{1}e^{-s\tau_{0}}$, the derivative of (28) with respect to $p$ is:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\boldsymbol{\rm S}}_{T}$ (29)
@@
$\dot{\boldsymbol{\rm S}}_{T}=$ $(\dot{\boldsymbol{\rm A}}_{1}h_{p}h_{s}+\boldsymbol{\rm A}_{1}(\frac{\partial{h_{p}}}{\partial{s}}h_{s}+h_{p}\frac{\partial{h_{s}}}{\partial{s}}))e^{-s\tau_{0}}$ (30)
$-\tau_{0}h_{p}h_{s}\boldsymbol{\rm A}_{1}e^{-s\tau_{0}}\dot{s}$
$=$ $\boldsymbol{\rm S}_{TD}-\tau_{0}\boldsymbol{\rm S}_{T}\dot{s}$
@@
$(\boldsymbol{\rm E}+\tau_{0}\boldsymbol{\rm S}_{T})\boldsymbol{\rm\phi}\dot{s}+(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm S}_{T})\dot{\boldsymbol{\rm\phi}}$ $=$ (31)
$-(s\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\boldsymbol{\rm S}_{TD})\boldsymbol{\rm\phi}$
@@ III-E Numerical Integration
We note that numerical integration schemes inherently introduce discretization error, which depends on the step size. This error can be practically eliminated by adding a Newton-based corrector step that recomputes the solution to within a prescribed tolerance, at additional computational cost.
@@ IV Case Study
In this section, we apply the continuation-based eigenvalue tracking formulation of Section III to two power system models. First, we assess accuracy based on a modified version of the IEEE 39-bus system [16]. We then discuss scalability based on a large-scale model of the Irish transmission system.
@@
We consider a modified version of the IEEE 39-bus system, wherein the machines at buses 30, 34, 35, 36, 37 are replaced by distributed energy resources of equal capacity. Each DER is modeled with an inner current control loop and outer voltage and frequency regulation loops [17].
@@
We begin by demonstrating the proposed tracking method on the system’s frequency regulation (FR) mode, also referred to as common low-frequency mode in the literature [18, 19]. This is a system-wide mode originating from the dynamics of frequency control loops [14]. For the examined system, the natural frequency of the FR is initially $0.005$ Hz. To illustrate the process, the frequency control droop constant $\rm R_{DER}$ of the five DERs is varied from $0.05$ to $0.5$. A single delay of $\tau=0.01$ s is introduced in the frequency signal of the power system stabilizer (PSS) connected to the synchronous machine (SM) at bus 31. The tracked eigenvalue, initially at $0.0276+\jmath 0.0165$, follows the trajectory shown in Fig. 1(a). Tracking based on (18) accurately captures the trajectory and shows how the oscillatory behavior of the FR mode is progressively suppressed as the relative share of frequency control shifts from the DERs to the SMs. At the point where the complex pair coalesces on the real axis, the eigenvalue becomes defective with algebraic multiplicity $2$ and geometric multiplicity $1$. This corresponds to a simple quadratic fold [20, 14]. As $\rm R_{DER}$ is further increased, the defective eigenvalue splits into two distinct real eigenvalues. The tracking method inherently follows one of the two emerging branches, as determined by the evolution of (18). In this example, the eigenvalue on the left branch of the fold is traced, and Fig. 1(a) shows two additional folds occurring along this trajectory. In practice, when such splitting occurs, the process is reinitialized by recomputing the eigendecomposition of $\boldsymbol{\rm P}$ and selecting the eigenpair corresponding to the branch of interest.
@@
We next assess the method in the presence of multiple delays. To this end, we introduce delays to the input signals of the system’s automatic voltage regulators, PSSs and DER frequency and voltage controllers, yielding a total of 20 time-delayed variables. The resulting $\boldsymbol{\rm P}$ has dimension $558\times 558$. For each delay $\tau_{j}$, the spectral discretization in [15] uses $N$ Chebyshev nodes in the interval $\tau_{j},0]$, which increases the size of $\boldsymbol{\rm P}$ by $N+1$. Thus, considering $N=10$, the dimension of the approximated linear eigenvalue problem grows by $220$ compared to the delay-free model. We then trace how variations in the SM droop constants affect the eigenvalues corresponding to (i) the system’s FR mode and (ii) the inter-machine mode with the largest participation in the rotor speeds of the SMs at buses $32$ and $39$, initially located at $0.4425\pm\jmath 0.2052$. The calculated trajectories, shown in Fig. 2 confirm that the proposed tracking method accurately captures eigenvalue trajectories in the presence of multiple delays.
We next examine the tracking accuracy when the delay magnitude is treated as the continuation parameter, as described in Section III-C. The analysis focuses on the local electromechanical mode of the SM at bus 31. The delay $\tau_{\rm PSS}$ in the input signal of the corresponding PSS is varied from $0.01$ to $0.6$ s, producing the trajectory shown in Fig. 3(a). As $\tau_{\rm PSS}$ increases, the traced eigenvalue becomes progressively less damped and eventually crosses into instability for $\tau_{\rm PSS}>0.1$ s. Further increasing $\tau_{\rm PSS}$ beyond $0.28$ s reverses this trend and eventually leads the system to regain small-signal stability for $\tau_{\rm PSS}>0.5$ s. This behavior is further validated via a time-domain simulation, considering a three-phase fault at bus 6, cleared after 80 ms by opening line 5–6. The response of the center-of-inertia (COI) frequency is shown in Fig. 3(b) and confirms the conclusions drawn by the eigenvalue tracking.
(a) Tracing local electromechanical mode as delay is increased.
@@ IV-B Irish System
We next assess the scalability of the proposed eigenvalue tracking method using a 1,502-bus dynamic model of the All-Island Irish Transmission System (AIITS). The delay-free DAE model has 1,629 states and 9,897 algebraic variables. Its dimension, $11,526\times 11,526$, renders the use of standard, dense QZ-based eigensolvers impractical. We first examine the accuracy of (18) in tracing poorly damped electromechanical modes of the AIITS. The DDAE model includes 28 delayed variables, affecting voltage and frequency measurements of the system’s AVRs and PSSs. The delays, ranging from $0.009$ to $0.022$ s, increase the dimension of the approximated linear eigenvalue problem to $11,834\times 11,834$. The effect of raising the gains of the system’s PSSs from $1.5$ to $10$, is shown in Fig. 4. Although the examined eigenvalues are tightly clustered in the complex plane, the proposed approach efficiently traces all of them.
@@ V Conclusion
$\;\;\,\boldsymbol{\rm h}_{1}=$ $\;\dot{\boldsymbol{\rm A}}_{0}-s_{\rm r}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{r}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{i}$
$\;\;\,\boldsymbol{\rm h}_{2}=$ $\;s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{i}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{r}$
@@
$\boldsymbol{\rm h}_{1}=$ $\dot{\boldsymbol{\rm A}}_{0}-s_{\rm r}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{r}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{i}$
$\boldsymbol{\rm h}_{2}=$ $s_{\rm i}\dot{\boldsymbol{\rm E}}+\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{i}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{r}$
@@
We set $\boldsymbol{\rm S}_{T}^{r}=\Re\{\boldsymbol{\rm S}_{T}\}$, $\boldsymbol{\rm S}_{T}^{i}=\Im\{\boldsymbol{\rm S}_{T}\}$, $\boldsymbol{\rm S}_{TD}^{r}=\Re\{\boldsymbol{\rm S}_{TD}\}$ and $\boldsymbol{\rm S}_{TD}^{i}=\Im\{\boldsymbol{\rm S}_{TD}\}$. Splitting real and imaginary parts of (31) leads to a system in the form of (18), where in this case $\boldsymbol{\rm M}_{3}$ is the same as in (18) and:
$\boldsymbol{\rm h}$ $=\begin{bmatrix}(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm S}_{TD}^{r})\boldsymbol{\phi}_{\rm r}+(s_{\rm i}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm S}_{TD}^{i})\boldsymbol{\phi}_{\rm i}\\ -(s_{\rm i}\dot{\boldsymbol{\rm E}}-\boldsymbol{\rm S}_{TD}^{i})\boldsymbol{\phi}_{\rm r}+(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+\boldsymbol{\rm S}_{TD}^{r})\boldsymbol{\phi}_{\rm i}\\ 0\\ 0\end{bmatrix}$
$\boldsymbol{\rm M}_{1}$ $=\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm S}_{T}^{r}&-s_{\rm i}\boldsymbol{\rm E}-\boldsymbol{\rm S}_{T}^{i}\\ s_{\rm i}\boldsymbol{\rm E}+\boldsymbol{\rm S}_{T}^{i}&s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-\boldsymbol{\rm S}_{T}^{r}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}$ $=\begin{bmatrix}\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\tau_{0}(\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm i})&\!\!-\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!-\!\tau_{0}(\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm i})\\ \boldsymbol{\rm E}\boldsymbol{\phi}_{\rm i}\!+\!\tau_{0}(\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm r}\!+\!\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm i})&\!\!\boldsymbol{\rm E}\boldsymbol{\phi}_{\rm r}\!+\!\tau_{0}(\boldsymbol{\rm S}_{T}^{r}\boldsymbol{\phi}_{\rm r}\!-\!\boldsymbol{\rm S}_{T}^{i}\boldsymbol{\phi}_{\rm i})\end{bmatrix}$

### v2

@@ III-D Communication Delays with Noise and Data Dropouts
The derivations above consider constant delays. In this section, we show how they can be generalized to capture the effects of realistic wide area measurement system (WAMS) latencies with noise and data packet dropouts. To this end, we consider the composite time-varying delay model proposed in [9], in which, for the purpose of SSSA, packet dropouts and noise are captured through the functions $h_{p}(s)$ and $h_{s}(s)$, respectively:
@@
where $\alpha$ and $b$ are the scale and shape factors of the Gamma distribution; $p_{dr}$ is the packet dropout rate; and $T$ is the normal delivery period for each data packet.
@@
where $\tau_{0}$ is the constant component of the WAMS delay. By setting ${\rm S}_{T}=h_{p}(s)h_{s}(s)e^{-s\tau_{0}}$, the derivative of (28) with respect to $p$ is:
$\dot{\boldsymbol{\rm P}}$ $=\dot{s}\boldsymbol{\rm E}+{s}\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\dot{\rm S}_{T}\boldsymbol{\rm A}_{1}-{\rm S}_{T}\dot{\boldsymbol{\rm A}}_{1}$ (29)
@@
$\dot{\rm S}_{T}=$ $(\frac{\partial{h_{p}}}{\partial{s}}h_{s}+h_{p}\frac{\partial{h_{s}}}{\partial{s}}-\tau_{0}h_{p}h_{s})e^{-s\tau_{0}}\dot{s}$ (30)
$=$ ${\rm S}_{TD}\dot{s}$
@@
$(s\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-{\rm S}_{T}\boldsymbol{\rm A}_{1})\dot{\boldsymbol{\rm\phi}}+(\boldsymbol{\rm E}-{\rm S}_{TD}\boldsymbol{\rm A}_{1})\boldsymbol{\rm\phi}\dot{s}$ $=$ (31)
$-(s\dot{\boldsymbol{\rm E}}-\dot{\boldsymbol{\rm A}}_{0}-\boldsymbol{\rm S}_{T}\dot{\boldsymbol{\rm A}}_{1})\boldsymbol{\rm\phi}$
@@ III-E Numerical Integration
We note that numerical integration schemes inherently introduce discretization error, which depends on the step size. This error can be practically eliminated by adding a Newton-based corrector step that recomputes the solution within a prescribed tolerance, at additional computational cost.
@@ IV Case Study
In this section, we apply the continuation-based eigenvalue tracking formulation of Section III to two power system models. First, we consider a modified version of the IEEE 39-bus system [4], and then a real-world-scale model of the Irish transmission system.
@@
We consider a modified version of the IEEE 39-bus system, wherein the machines at buses 30, 34, 35, 36, 37 are replaced by distributed energy resources (DERs) of equal capacity. Each distributed energy resource (DER) is modeled with an inner current control loop and outer voltage and frequency regulation loops [13].
@@
We begin by demonstrating the proposed tracking method on the system’s frequency regulation (FR) mode, also referred to as common low-frequency mode in the literature [19, 11]. This is a system-wide mode originating from the dynamics of frequency control loops [1]. For the examined system, the natural frequency of the FR mode is initially $0.005$ Hz. To illustrate the process, the frequency control droop constant $\rm R_{DER}$ of the five DERs is varied from $0.05$ to $0.5$. A single delay of $\tau=0.01$ s is introduced in the frequency signal of the power system stabilizer (PSS) connected to the synchronous machine (SM) at bus 31. The traced eigenvalue, initially at $0.0276+\jmath 0.0165$, follows the trajectory shown in Fig. 1(a). Tracking based on (18) accurately captures the trajectory and shows how the oscillatory behavior of the FR mode is progressively suppressed as the relative share of frequency control shifts from the DERs to the SMs. At the point where the complex pair coalesces on the real axis, the eigenvalue becomes defective with algebraic multiplicity $2$ and geometric multiplicity $1$. This corresponds to a simple quadratic fold [15, 1]. As $\rm R_{DER}$ is further increased, the defective eigenvalue splits into two distinct real eigenvalues. The tracking method inherently follows one of the two emerging branches, as determined by the evolution of (18). In this example, the eigenvalue on the left branch of the fold is traced, and Fig. 1(a) shows two additional folds occurring along this trajectory. In practice, when such a splitting occurs, the process is reinitialized by recomputing the eigendecomposition of $\boldsymbol{\rm P}$ and selecting the eigenpair corresponding to the branch of interest.
@@
We next assess the method in the presence of multiple delays. To this end, we introduce delays to the input signals of the system’s automatic voltage regulators, PSSs and DER frequency and voltage controllers, yielding a total of 20 time-delayed variables. The resulting $\boldsymbol{\rm P}$ has dimension $\times558$. For each delay $\tau_{j}$, the spectral discretization in [6] uses $N$ Chebyshev nodes in the interval $\tau_{j},0]$, which increases the size of $\boldsymbol{\rm P}$ by $N+1$. Thus, considering $N=10$, the dimension of the approximated linear eigenvalue problem grows by $220$ compared to the delay-free model. We then trace how variations in the SM droop constants affect the eigenvalues corresponding to (i) the system’s FR mode and (ii) the inter-machine mode with the largest participation in the rotor speeds of the SMs at buses $32$ and $39$, initially located at $0.4425\pm\jmath 0.2052$. The calculated trajectories, shown in Fig. 2, confirm that the proposed tracking method accurately captures eigenvalue trajectories in the presence of multiple delays.
We next examine the tracking accuracy when the delay magnitude is treated as the continuation parameter, as described in Section III-C. The analysis focuses on the local electromechanical mode of the SM at bus 31. The delay $\tau_{\rm PSS}$ in the input signal of the corresponding PSS is varied from $0.01$ to $0.6$ s, producing the trajectory shown in Fig. 3(a). As $\tau_{\rm PSS}$ increases, the traced eigenvalue becomes progressively less damped and eventually crosses into instability for $\tau_{\rm PSS}>0.1$ s. Further increasing $\tau_{\rm PSS}$ beyond $0.28$ s reverses this trend and eventually leads the system to regain small-signal stability for $\tau_{\rm PSS}>0.5$ s. This behavior is further validated via a time-domain simulation, considering a three-phase fault at bus 6, cleared after 80 ms by opening line 5–6. The response of the corresponding SM rotor speed is shown in Fig. 3(b) and confirms the conclusions drawn by the eigenvalue tracking. In the case of $\tau_{\rm PSS}=0.3$ s, the PSS’s output introduces a destabilizing negative damping component to the SM’s electrical torque, causing the rotor speed to diverge.
(a) Tracing a local electromechanical mode as delay is increased.
@@ IV-B Irish System
We assess the scalability of the proposed eigenvalue tracking method using a 1,502-bus dynamic model of the All-Island Irish Transmission System (AIITS). The delay-free DAE model has 1,629 states and 9,897 algebraic variables. Its dimension, $\times11,526$, renders the use of standard, dense QZ-based eigensolvers impractical. We first examine the accuracy of (18) in tracing poorly damped electromechanical modes of the AIITS. The DDAE model includes 28 delayed variables, affecting voltage and frequency measurements of the system’s AVRs and PSSs. The delays, ranging from $0.009$ to $0.022$ s, increase the dimension of the approximated linear eigenvalue problem to $\times11,834$. The effect of raising the gains of the system’s PSSs from $1.5$ to $10$, is shown in Fig. 4. Although the examined eigenvalues are tightly clustered in the complex plane, the proposed approach efficiently traces all of them.
@@ V Conclusion
$\;\;\,\boldsymbol{\rm h}_{1}=$ $\;-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{r}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{i}$
$\;\;\,\boldsymbol{\rm h}_{2}=$ $\;\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{i}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{r}$
@@
$\boldsymbol{\rm h}_{1}=$ $-\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{r}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{i}$
$\boldsymbol{\rm h}_{2}=$ $\boldsymbol{\rm A}_{\ell}h^{\ell}_{r}s_{i}-\boldsymbol{\rm A}_{\ell}h^{\ell}_{i}s_{r}$
@@
We set ${\rm S}_{T}^{r}=\Re\{{\rm S}_{T}\}$, ${\rm S}_{T}^{i}=\Im\{{\rm S}_{T}\}$, $\tilde{\boldsymbol{\rm S}}_{TD}^{r}=\Re\{{\rm S}_{TD}\}\boldsymbol{\rm A}_{1}$ and $\tilde{\boldsymbol{\rm S}}_{TD}^{i}=\Im\{{\rm S}_{TD}\}\boldsymbol{\rm A}_{1}$. Splitting real and imaginary parts of (31) leads to a system in the form of (18), where in this case $\boldsymbol{\rm M}_{3}$ is the same as in (18) and:
$\boldsymbol{\rm h}$ $=\begin{bmatrix}(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+{\rm S}_{T}^{r}\dot{\boldsymbol{\rm A}}_{1})\boldsymbol{\phi}_{\rm r}+(s_{\rm i}\dot{\boldsymbol{\rm E}}-{\rm S}_{T}^{i}\dot{\boldsymbol{\rm A}}_{1})\boldsymbol{\phi}_{\rm i}\\ -(s_{\rm i}\dot{\boldsymbol{\rm E}}-{\rm S}_{T}^{i}\dot{\boldsymbol{\rm A}}_{1})\boldsymbol{\phi}_{\rm r}+(-s_{\rm r}\dot{\boldsymbol{\rm E}}+\dot{\boldsymbol{\rm A}}_{0}+{\rm S}_{T}^{r}\dot{\boldsymbol{\rm A}}_{1})\boldsymbol{\phi}_{\rm i}\\ 0\\ 0\end{bmatrix}$
$\boldsymbol{\rm M}_{1}$ $=\begin{bmatrix}s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-{\rm S}_{T}^{r}\boldsymbol{\rm A}_{1}&-s_{\rm i}\boldsymbol{\rm E}+{\rm S}_{T}^{i}\boldsymbol{\rm A}_{1}\\ s_{\rm i}\boldsymbol{\rm E}-{\rm S}_{T}^{i}\boldsymbol{\rm A}_{1}&s_{\rm r}\boldsymbol{\rm E}-\boldsymbol{\rm A}_{0}-{\rm S}_{T}^{r}\boldsymbol{\rm A}_{1}\end{bmatrix}$
$\boldsymbol{\rm M}_{2}$ $=\begin{bmatrix}(\boldsymbol{\rm E}\!-\tilde{\boldsymbol{\rm S}}_{TD}^{r})\boldsymbol{\phi}_{\rm r}\!+\!\tilde{\boldsymbol{\rm S}}_{TD}^{i}\boldsymbol{\phi}_{\rm i}&\!\!-(\boldsymbol{\rm E}\!-\!\tilde{\boldsymbol{\rm S}}_{TD}^{r})\boldsymbol{\phi}_{\rm i}\!+\!\tilde{\boldsymbol{\rm S}}_{TD}^{i}\boldsymbol{\phi}_{\rm r}\\ (\boldsymbol{\rm E}\!-\!\tilde{\boldsymbol{\rm S}}_{TD}^{r})\boldsymbol{\phi}_{\rm i}\!-\!\tilde{\boldsymbol{\rm S}}_{TD}^{i}\boldsymbol{\phi}_{\rm r}&\!\!(\boldsymbol{\rm E}\!-\tilde{\boldsymbol{\rm S}}_{TD}^{r})\boldsymbol{\phi}_{\rm r}\!+\!\tilde{\boldsymbol{\rm S}}_{TD}^{i}\boldsymbol{\phi}_{\rm i}\end{bmatrix}$